OpenStudy (anonymous):

Please Help! The expression sinx (cosx cotx - sinx) simplifies to_______? 1) cot^2x-sin^2x 2) 1 3) cos(2x)

OpenStudy (anonymous):

3 is correct cos*2x-sin*2x=cos2x

OpenStudy (vishweshshrimali5):

\[cosx cotx - sinx = cosx * \cos x/\sin x - \sin x = (\cos ^2 x - \sin^2x)/\sin x = \cos 2x/\sin x\] \[answer = \cos 2x/\sin x*\sin x = \cos 2x\]